Calculate the probability of drawing four cards out of a standard deck.
P(4 of a kind)
P(All the same suit)
P(All the same #)
Thanks So much!

Answer by stanbon(75887)

Calculate the probability of drawing four cards out of a standard deck.
P(4 of a kind)
There are 13 "kinds" so 13 ways to succeed
There are 52C4 different 4-card hands
P(4 of a kind) = 13/52C4 = 0.000048...
----------------------------
P(All the same suit)
There are 4 ways to pick a suit
There are 13C4 4-card sets in each suit
P(All the same suit) = [4*13C4]/52C4 = 2860/270725 = 0.0106...
------------------------
P(All the same #)
I'm going to assume "#" does not include Ace,King,Queen,Jack
There are 9 ways to pick a number.
There 1 way to get all four cards of that number
P(All the same #) = 9/52C4 = 0.000033244...
Cheers,
Stan H.
--